let n€N consider the set of nxn symmetric matrices over R with the usual addition and multiplication by a scalar.

a) Show that this set with the given operations is a vector subspace of Mnn

b) What is the dimension of this vector subspace?

c) Find a basis for the vector space of 2x2 symmetric matrices.

1
Expert's answer
2021-11-03T18:38:13-0400

a)

We have


"V=M_n(\R)"


is the vector space over "\R" .


Now,

denote "A_{ij}" is the matrix whose "(i,j)^{th}" entry is "1" if "i=j"

and ;if



"i\ne j,\,\forall i,j\in \{1,\dots,n\}"

Also denote the collection of all such matrix by "X" ,

thus;


"W=\text{span}(X)" is the smallest subspace


such that "I_{n\times n}\in W" .


Define



"M_n^s(\R):=\{A\in M_n(\R):A=A^T\}"



Clearly, "M_n^s(\R)" is the set of all symmetric matrix. we show that it is subspace of "M_n(\R)" .


Let, "a,b\in \R" and "A,B\in M_n^s(\R)" , thus "A^T=A,B^T=B"


Now, consider the liner combination "aA+bB" .


Note that



"(aA+bB)^T=(bB)^T+(aA)^T\\
=(aA)^T+(bB)^T\\
=aA^T+bB^T\\
=aA+bB"

Thus,we can see that



"aA+bB\in M_n^s(\R)"



Hence,it can be seen that



"M_n^s(\R)" is subspace of "M_n(\R)"


b)


Let

"A=\\begin{bmatrix} \n\t\t a_{11} & a_{12}\\\\ \n\t\t a_{21}& a_{22} \n\t\t\\end{bmatrix}"

be an arbitrary element in the subspace W.

Then since

AT=A, we have

"\\begin{bmatrix} \n\t\t a_{11} & a_{21}\\\\ \n\t\t a_{12}& a_{22} \n\t\t\\end{bmatrix}=\\begin{bmatrix} \n\t\t a_{11} & a_{12}\\\\ \n\t\t a_{21}& a_{22} \n\t\t\\end{bmatrix}."

This implies that


"a_{12}=a_{21}" , and hence


"A=\n\\begin{bmatrix} \n\t\t a_{11} & a_{12}\\\\ \n\t\t a_{12}& a_{22} \n\t\t\\end{bmatrix}\\\\"


"=a_{11}\\begin{bmatrix} \n\t\t 1 & 0\\\\ \n\t\t 0& 0 \n\t\t\\end{bmatrix}+a_{12}\\begin{bmatrix} \n\t\t 0 & 1\\\\ \n\t\t 1& 0 \n\t\t\\end{bmatrix}+a_{22}\\begin{bmatrix} \n\t\t 0 & 0\\\\ \n\t\t 0& 1 \n\t\t\\end{bmatrix}."

Let

B={v1,v2,v3}, where 

v1,v2,v3 are 2×2

 matrices appearing in the above linear combination of A.

Note that these matrices are symmetric.

Hence we showed that any element in W

W is a linear combination of matrices in B.

Thus Bi s a spanning set for the subspace W.

We show that B is linearly independent.

Suppose that we have

"c_1v_1+c_2v_2+c_3v_3=\\begin{bmatrix} \n\t\t 0 & 0\\\\ \n\t\t 0& 0 \n\t\t\\end{bmatrix}."

Then it follows that

["\\begin{bmatrix} \n\t\t c_1 & c_2\\\\ \n\t\t c_2& c_3 \n\t\t\\end{bmatrix}=\\begin{bmatrix} \n\t\t 0 & 0\\\\ \n\t\t 0& 0 \n\t\t\\end{bmatrix}."

Thus

c1=c2=c3=and the setB

is linearly independent.As 

B is a linearly independent spanning set, we conclude that 

B is a basis for the subspace 

W.


c)

Recall that the dimension of a subspace is the number of vectors in a basis of the subspace.

In part (b), we found that 

B={v1,v2,v3} is a basis for the subspace W

W.

As 

B consists of three vectors, the dimension of W

W is 3.